# Hamiltonian reductions as affine closures of cotangent bundles

### v1

@@ 3.3. Hamiltonian reduction
so the symplectic form is independent of the choice of the Hamiltonian slice $U$. ∎
@@
By Lemma 2.6, the shell $N$ is normal and hence so is N//GN/\!/G and $N=N^{\operatorname{red}}$. In particular, since $N$ has FPIG by assumption, the open subset $N_{\operatorname{pr}}$ consists precisely of the closed principal orbits by Luna’s slice theorem. This implies
@@
Now let $(H)$ be a closed non-principal isotropy class. By Lemmas 3.4, 3.5 and the assumption, we have
@@
This implies that $(N/\!/G)\setminus(N/\!/G)_{\operatorname{pr}}$ has codimension at least $4$, and the canonical symplectic form on $(N/\!/G)_{\operatorname{pr}}$ obtained in Lemma 3.6 then extends to a $2$-form on $(N/\!/G)_{\operatorname{reg}}$, which is again symplectic because its degenerate locus has codimension one if it is non-empty. Finally, since $(N/\!/G)\setminus(N/\!/G)_{\operatorname{reg}}\subset(N/\!/G)\setminus(N/\!/G)_{\operatorname{pr}}$ has codmension at least $4$, it follows that X///G=N//GX/\!\!/\!\!/G=N/\!/G is a symplectic variety by Flenner’s theorem [Fle88] and with terminal singularities by [Nam01]. ∎
@@
A modified version of Proposition 3.7 also holds for Cohen–Macaulay affine quasi-Hamiltonian $G$-varieties. The main difficulty in the singular setting is that the Hamiltonian slice theorem does not hold at singular points and we have no control on the dimension of $(N/\!/G)_{(H)}$ at closed orbits lying in the locus $N\cap X_{\operatorname{sing}}$. However, if the image $\pi_{N}(N\cap X_{\operatorname{sing}})$ has codimension at least $4$ in N//GN/\!/G (for example, when $\dim X_{\operatorname{sing}}\leq\dim X-2\dim G-4$), then Proposition 3.7 remains valid when we restrict to points $x\in(N^{\operatorname{red}}\cap X_{\operatorname{reg}})\setminus N_{\operatorname{pr}}^{\operatorname{red}}$. We leave the verification of these details to the interested reader.
@@ 4. Hamiltonian reduction of cotangent bundles
We now apply Proposition 3.7 to study the Hamiltonian reduction of the cotangent bundle of a smooth affine variety with a $2$-large $G$-action, thereby proving Theorems 1.1 and 1.2.
Throughout this section, we fix the following notation. Given an irreducible non-singular affine $G$-variety $Y$, we denote by $\mu\colon T^{*}Y\rightarrow\mathfrak{g}^{*}$ the canonical moment map. The shell of $\mu$ is denoted by $N$. The GIT quotient of $Y$ under $G$ is written as $Z\coloneqq Y/\!/G$, with quotient map $\pi_{Y}\colon Y\rightarrow Z$. The Hamiltonian reduction of $T^{*}Y$ is denoted as $M\coloneqq T^{*}Y/\!\!/\!\!/G=N/\!/G$ with the corresponding quotient $\pi_{N}\colon N\rightarrow M$.
@@ 4.1. Large actions
is a codimension $d$ linear subspace of $T^{*}_{y}Y$. In particular, as $Y_{(0)}$ is irreducible, the variety $N^{0}$ is also irreducible. Now we observe
@@
First we note that $N$ is an irreducible normal variety by Proposition 4.2. Moreover, since $Y$ has FPIG, to prove (4.3.1), it suffices to show that any point $x\in p^{-1}(Y_{\operatorname{pr}})\cap N$ has closed $G$-orbit such that $G_{x}=G_{y}$.
Let $x\in p^{-1}(Y_{\operatorname{pr}})\cap N$ with $y\coloneqq p(x)\in Y_{\operatorname{pr}}$. Moreover, since $Y$ has FPIG, the open subset $Y_{\operatorname{pr}}$ consists of closed principal orbits, so $O_{y}=Gy$ is closed in $Y$ and $G_{x}\subset G_{y}$ is a finite subgroup. Moreover, the orbit $O_{x}=Gx$ is contained in the closed subset
@@
of $X$ such that $O_{x}\cong G\times_{G_{y}}O^{\prime}_{x}$, where $G_{y}$ acts on $T^{*}_{y}Y$ by the isotropy action and the orbit $O^{\prime}_{x}=G_{y}x\subset T_{y}^{*}Y$ is a finite set. So $O_{x}$ is closed in $p^{-1}(O_{y})$ and hence is closed in $T^{*}Y$.
@@
Now note that $Z_{\operatorname{pr}}=Z_{\operatorname{reg}}$ by [Sch95, Theorem 9.2], so it remains to prove that $\operatorname{codim}M\setminus M^{\diamond}\geq 2$. Let $N^{r}\coloneqq p^{-1}(Y_{(r)})\cap N$ for any $0\leq r\leq\dim G$. Then, for $1\leq r\leq\dim G$, by (4.2.1), we have
@@
which implies $\dim(M\setminus M^{\diamond})\leq\dim M-2$, and hence (1.1.2) follows. ∎
@@
Let $Y$ be an irreducible non-singular affine $G$-variety with FIPG. Assume that the identity component $G^{0}$ is a torus. By [Sch95, Proposition 10.1], $Y$ is $k$-large if and only if $Y$ is $k$-principal. If $G$ acts on $Y$ freely outside a codimension at least two locus, then $Y$ is $2$-large. Theorem 1.1 implies that $T^{*}Y/\!\!/\!\!/G$ is a symplectic variety with terminal singularities which is the affine closure of $T^{*}(Y/\!/G)_{\operatorname{pr}}$. One particular example is when $G$ is a finite group acting freely on $Y^{\diamond}\subset Y$ with $\operatorname{codim}Y\setminus Y^{\diamond}\geq 2$, then $T^{*}Y/G$ is the affine closure of $T^{*}(Y^{\diamond}/G)$, which was previously remarked in [FL25a, Example 3.4].
@@ 4.2. Factoriality
For an irreducible non-singular affine $G$-variety $Y$, we remark that it follows from standard properties of affine GIT that the natural projections $Y\rightarrow Z$ and $Y_{\operatorname{pr}}\rightarrow Z_{\operatorname{pr}}$ are both quotients by $G$ in the sense of [KKV89, 4.1].
@@
We have $Z_{\operatorname{pr}}=Z_{\operatorname{reg}}$ by [Sch95, Theorem 9.12]. In particular, since $Z$ is normal, we have $\operatorname{Cl}(Z)\cong\operatorname{Pic}(Z_{\operatorname{pr}})$. On the other hand, since $Y$ is $2$-principal, then the open subset $Y_{\operatorname{pr}}$ is $G$-stable with $\operatorname{codim}(Y\setminus Y_{\operatorname{pr}})\geq 2$. In particular, we have
@@
By [KKV89, Lemma 2.2], we have an exact sequence
@@
Then applying [KKV89, Proposition 4.2] to the quotient $Y_{\operatorname{pr}}\rightarrow Z_{\operatorname{pr}}$ yields that $\operatorname{Pic}(Z_{\operatorname{pr}})$ is a subgroup of $\operatorname{Pic}_{G}(Y_{\operatorname{pr}})$, so the result follows. ∎
@@
implies that $H^{1}_{\operatorname{alg}}(G,\mathcal{O}(Y)^{\times})$ is trivial (resp. finite) if the character group $\mathfrak{X}(G)$ of $G$ is trivial (resp. finite) and hence if $G_{\operatorname{ab}}$ is trivial (resp. finite).
Finally, recall that the singular locus of terminal singularities has codimension at least three, so the result follows from [Fu03, Corollary 1.3], which shows that a locally $\mathbb{Q}$-factorial symplectic variety does not admit any symplectic resolution if its singular locus is not of pure codimension two. ∎
@@
Let $Y$ be an irreducible non-singular affine $G$-variety which is $2$-large. Assume that $\mathcal{O}(Y)^{\times}=\mathbb{C}^{\times}$, $Y^{G}\not$$\emptyset$ and $\operatorname{Pic}(Y)=0$.
@@
    The divisor class group $\operatorname{Cl}(Z)$ is isomorphic to the kernel of $\mathfrak{X}(G)\rightarrow\mathfrak{X}(H)$, where $H$ is a principal isotropy group.
@@
Applying [KKV89, Corollary 5.3] to the quotient $Y\rightarrow Z$ yields $\operatorname{Pic}(Z)=0$. Since $M$ is the affine closure of $T^{*}Z_{\operatorname{pr}}$ by Theorem 1.1, there exists a natural projection $p_{Z}\colon M\rightarrow Z$ such that the restriction $p_{Z}|_{T^{*}Z_{\operatorname{pr}}}\colon T^{*}Z_{\operatorname{pr}}\rightarrow Z_{\operatorname{pr}}$ identifies to the natural projection. Let $\mathbf{Z}$ be the zero section of $p_{Z}$; that is, the closure of the zero section $\mathbf{Z}_{\operatorname{pr}}$ of $T^{*}Z_{\operatorname{pr}}\rightarrow Z_{\operatorname{pr}}$ in $M$ or equivalently the image $\pi_{N}(\mathbf{Y})$, where $\mathbf{Y}$ $(\subset N)$ is the zero section of $T^{*}Y\rightarrow Y$. Then the natural morphism
$\iota=p_{Z}|_{\mathbf{Z}}\colon\mathbf{Z}\longrightarrow Z$
is an isomorphism. Consider the following commutative diagram:
@@
By Theorem 1.2 and Proposition 4.6, we only need to prove that if $M$ is $\mathbb{Q}$-factorial, then $G_{\operatorname{ab}}$ is finite. Assume to the contrary that $G_{\operatorname{ab}}$ is not finite. Then we can choose a surjective character $\chi\colon G\rightarrow\mathbb{C}^{\times}$. In particular, $\chi^{n}\not=1$ for all $n\geq 1$. Nevertheless, as the principal isotropy gruop $H$ is finite, there exists $m\in\mathbb{Z}_{>0}$ such that $\chi^{m}|_{H}=1$, which then implies that $\operatorname{Cl}(Z)$ is not torsion and hence $M$ is not $\mathbb{Q}$-factorial by Proposition 4.6, which is absurd. ∎
@@ 4.3. Examples
We collect some examples to illustrate the sharpness of Theorems 1.1 and 1.2.
@@
Let $V=(\mathfrak{sl}_{2})^{2}$, as a representation of $G=\operatorname{PGL}_{2}$, which is $1$-large, but not $2$-large. Then the Hamiltonian reduction $T^{*}V/\!\!/\!\!/G$ identifies with the locus of square zero matrices in $\mathfrak{sp}_{4}$, and in particular is a symplectic variety. Nevertheless, its singular locus is the codimension two locus consisting of rank one matrices in $\mathfrak{sp}_{4}$ and hence it is not terminal, so the part (1.1.1) of Theorem 1.1 is optimal.
@@
It is easy to see that for any $k$, the Hamiltonian reduction $T^{*}V_{k}/\!\!/\!\!/G$ is isomorphic to the minimal nilpotent orbit closure $\overline{\mathbf{O}}_{{\rm min}}$ in $\mathfrak{sl}_{n}$. For $z\in V_{k}$, the $G$-orbit GzGz is closed in $V_{k}$ if and only if $z=0$ or $z\in V_{k}\setminus(F_{k}\cup F^{\prime}_{n-k})$, where $F_{k}=\{x_{1}=\cdots=x_{k}=0\}$ and $F^{\prime}_{n-k}=\{x_{k+1}=\cdots=x_{n}=0\}$. It follows that $(V_{k})_{\operatorname{pr}}=V_{k}\setminus(F_{k}\cup F^{\prime}_{n-k})$, thus $V_{k}$ is 2-large if and only if $2\leq k\leq n-2$. In this case, Theorem 1.1 implies that $\overline{\mathbf{O}}_{{\rm min}}$ is the affine closure of $T^{*}((V_{k}/\!/G)_{\operatorname{pr}})$. On the other hand, $(V_{k}/\!/G)_{\operatorname{pr}}$ is just the punctured affine cone of the Segre variety $\mathbb{P}^{k-1}\times\mathbb{P}^{n-k-1}\subset\mathbb{P}^{k(n-k)-1}$. This shows that $\overline{\mathbf{O}}_{{\rm min}}$ is the affine closure of the cotangent bundle of the punctured affine cone of a Segre variety, which is a special case of [FL25a, Theorem 1.5] but with a new proof. For $k=0$, $V_{0}/\!/G$ is just one point. For $k=1$, $V_{1}$ is $1$-large with $V_{1}/\!/G\simeq\mathbb{C}^{n-1}$ and $(V_{1}/\!/G)_{\operatorname{pr}}\simeq\mathbb{C}^{n-1}\setminus\{0\}$. It follows that the affine closure of $T^{*}(V_{1}/\!/G)_{\operatorname{pr}}$ is $T^{*}\mathbb{C}^{n-1}$ which is different from $\overline{\mathbf{O}}_{{\rm min}}$. Hence the part (1.1.2) of Theorem 1.1 is sharp.
@@
Following the notation in Example 4.9, then we have $G_{\operatorname{ab}}=G=\mathbb{C}^{\times}$. Moreover, it is well-known that $\overline{\mathbf{O}}_{{\rm min}}$ is not $\mathbb{Q}$-factorial and admits a symplectic resolution by collapsing the zero section of $T^{*}\mathbb{P}^{n-1}$.
Now let $n=2k$ and let $G^{\prime}\coloneqq\mathbb{C}^{\times}\rtimes\mathbb{Z}_{2}$, where $s\in\mathbb{Z}_{2}$ acts on $\mathbb{C}^{\times}$ by $s(\lambda)=\lambda^{-1}$. Then $G^{\prime},G^{\prime}]=\mathbb{C}^{\times}$ and $G_{\operatorname{ab}}\cong\mathbb{Z}_{2}$. Let $s$ act on $V_{k}=\mathbb{C}^{2k}$ as follows:
@@
The nilpotent closure $\overline{\textbf{O}}_{[2^{2},1^{2k-4}]}$ is the next to the minimal orbit closure in the sense that it is contained in all other nilpotent orbit closures, and $\operatorname{Cl}(\overline{\textbf{O}}_{[2^{2},1^{2k-4}]})\cong\mathbb{Z}_{2}$ by [Fu03, Corollary 2.7]. These examples show that we cannot remove the triviality and finiteness conditions on $G_{\operatorname{ab}}$ in Theorem 1.2. In this sense, our Theorem 1.2 is also optimal.

### v2

@@ 3.3. Hamiltonian reduction
so, the symplectic form is independent of the choice of the Hamiltonian slice $U$. ∎
@@
By Lemma 2.6, the shell $N$ is normal, so $N=N^{\operatorname{red}}$ and consequently N//GN/\!/G is also normal. In particular, since $N$ has FPIG by assumption, the open subset $N_{\operatorname{pr}}$ consists precisely of the closed principal orbits by Luna’s slice theorem. This implies
@@
Now, let $(H)$ be a closed non-principal isotropy class. By Lemmas 3.4, 3.5 and the assumption, we have
@@
This implies that $(N/\!/G)\setminus(N/\!/G)_{\operatorname{pr}}$ has codimension at least four, and the canonical symplectic form on $(N/\!/G)_{\operatorname{pr}}$ obtained in Lemma 3.6 then extends to a $2$-form on $(N/\!/G)_{\operatorname{reg}}$, which is again symplectic because its degenerate locus has codimension one if it is non-empty. Finally, since $(N/\!/G)\setminus(N/\!/G)_{\operatorname{reg}}\subset(N/\!/G)\setminus(N/\!/G)_{\operatorname{pr}}$ has codimension at least four, it follows that X///G=N//GX/\!\!/\!\!/G=N/\!/G is a symplectic variety by Flenner’s theorem [Fle88] and with terminal singularities by [Nam01]. ∎
@@
A modified version of Proposition 3.7 also holds for Cohen–Macaulay affine quasi-Hamiltonian $G$-varieties. The main difficulty in the singular setting is that the Hamiltonian slice theorem does not hold at singular points, and we have no control on the dimension of $(N/\!/G)_{(H)}$ at closed orbits lying in the locus $N\cap X_{\operatorname{sing}}$. However, if the image $\pi_{N}(N\cap X_{\operatorname{sing}})$ has codimension at least four in N//GN/\!/G (for example, when $\dim X_{\operatorname{sing}}\leq\dim X-2\dim G-4$), then Proposition 3.7 remains valid when we restrict it to points $x\in(N^{\operatorname{red}}\cap X_{\operatorname{reg}})\setminus N_{\operatorname{pr}}^{\operatorname{red}}$. We leave the verification of these details to the interested reader.
@@ 4. Hamiltonian reduction of cotangent bundles
We now apply Proposition 3.7 to study the Hamiltonian reduction of the cotangent bundle of a smooth affine variety with a $2$-large $G$-action, thus proving Theorems 1.1 and 1.3.
Throughout this section, we fix the following notation. Given an irreducible non-singular affine $G$-variety $Y$, we denote by $\mu\colon T^{*}Y\rightarrow\mathfrak{g}^{*}$ the canonical moment map. The shell of $\mu$ is denoted by $N$. The GIT quotient of $Y$ under $G$ is written as $Z\coloneqq Y/\!/G$, with the quotient map $\pi_{Y}\colon Y\rightarrow Z$. The Hamiltonian reduction of $T^{*}Y$ is denoted as $M\coloneqq T^{*}Y/\!\!/\!\!/G=N/\!/G$ with the corresponding quotient $\pi_{N}\colon N\rightarrow M$.
@@ 4.1. Large actions
is a codimension $d$ linear subspace of $T^{*}_{y}Y$. In particular, since $Y_{(0)}$ is irreducible, the variety $N^{0}$ is also irreducible. Now we observe
@@
Firstly, we note that $N$ is an irreducible normal variety by Proposition 4.2. Moreover, since $Y$ has FPIG, to prove (4.3.1), it suffices to show that any point $x\in p^{-1}(Y_{\operatorname{pr}})\cap N$ has a closed $G$-orbit such that $G_{x}=G_{y}$.
Let $x\in p^{-1}(Y_{\operatorname{pr}})\cap N$ and $y\coloneqq p(x)\in Y_{\operatorname{pr}}$. Moreover, since $Y$ has FPIG, the open subset $Y_{\operatorname{pr}}$ consists of closed principal orbits, so $O_{y}=Gy$ is closed in $Y$ and $G_{x}\subset G_{y}$ is a finite subgroup. Moreover, the orbit $O_{x}=Gx$ is contained in the closed subset
@@
of $X$ such that $O_{x}\cong G\times_{G_{y}}O^{\prime}_{x}$, where $G_{y}$ acts on $T^{*}_{y}Y$ by the isotropy action and the orbit $O^{\prime}_{x}=G_{y}x\subset T_{y}^{*}Y$ is a finite set. So $O_{x}$ is closed in $p^{-1}(O_{y})$ and is therefore closed in $T^{*}Y$.
@@
Since $Y$ is $2$-principal, the closed subset $Y\setminus Y_{\operatorname{pr}}$ has codimension at least two in $Y$, so that $Z_{\operatorname{reg}}\setminus Z_{\operatorname{pr}}$ has codimension at least two in $Z_{\operatorname{reg}}$. This implies that the restriction $\mathcal{O}(T^{*}Z_{\operatorname{reg}})\rightarrow\mathcal{O}(T^{*}Z_{\operatorname{pr}})$ is an isomorphism of $\mathbb{C}$-algebras. So it remains to prove that $\operatorname{codim}M\setminus M^{\diamond}\geq 2$. Let $N^{r}\coloneqq p^{-1}(Y_{(r)})\cap N$ for any $0\leq r\leq\dim G$. Then, for $1\leq r\leq\dim G$, by (4.2.1), we have
@@
which implies $\dim(M\setminus M^{\diamond})\leq\dim M-2$, and hence (1.1.2) follows.
For (1.1.3), firstly note that if $Z$ is non-singular, then $T^{*}Z$ is an irreducible non-singular affine variety such that $M=T^{*}Z$, so $M$ is also non-singular.
For the converse, we assume that $Z$ is singular. By (1.1.2), there exists a natural projection $p_{Z}\colon M\rightarrow Z$ such that the restriction $p_{Z}|_{T^{*}Z_{\operatorname{pr}}}\colon T^{*}Z_{\operatorname{pr}}\rightarrow Z_{\operatorname{pr}}$ identifies to the natural projection. Let $\mathbf{Z}$ be the zero section of $p_{Z}$; that is, the closure of the zero section $\mathbf{Z}_{\operatorname{pr}}$ of $T^{*}Z_{\operatorname{pr}}\rightarrow Z_{\operatorname{pr}}$ in $M$ or equivalently $M^{\mathbb{C}^{*}}$, where the $\mathbb{C}^{*}$-action on $M$ is induced by the natural scaling on the fibers of $T^{*}Z_{\operatorname{pr}}\rightarrow Z_{\operatorname{pr}}$. Then the restriction $\iota=p_{Z}|_{\mathbf{Z}}\colon\mathbf{Z}\longrightarrow Z$ is an isomorphism.
Let $z$ be a singular point of $Z$ and let $\mathbf{z}\in\mathbf{Z}$ be the point such that $\iota(\mathbf{z})=z$. Let $F$ be an irreducible component of the fiber of $p_{Z}$ passing through $\mathbf{z}$. Then $\dim F\geq\dim Z$ by the upper semi-continuity. Let $T_{z}Z$, $T_{\mathbf{z}}M$ and $T_{\mathbf{z}}F$ be the Zariski tangent spaces. Then we have the following complex of $\mathbb{C}$-linear maps given by the tangent maps:
$T_{\mathbf{z}}F\longrightarrow T_{\mathbf{z}}M\longrightarrow T_{z}Z,$
where the first map is injective. Moreover, since $\iota\colon\mathbf{Z}\rightarrow Z$ is an isomorphism, the following composition
$T_{\mathbf{z}}\mathbf{Z}\longrightarrow T_{\mathbf{z}}M\longrightarrow T_{z}Z$
is an isomorphism, so the last map is surjective. On the other hand, since $z$ is a singular point of $Z$, we have $\dim T_{z}Z>\dim Z$, which implies
$\dim T_{\mathbf{z}}M\geq\dim T_{\mathbf{z}}F+\dim T_{z}Z>\dim F+\dim Z\geq 2\dim Z=\dim M.$
Hence $M$ is singular at $\mathbf{z}$ by Jacobian’s criterion. ∎
Remark 4.4.
For an irreducible non-singular affine $G$-variety $Y$ that is $2$-large, we have $Z_{\operatorname{pr}}=Z_{\operatorname{reg}}$ by [Sch95, Theorem 9.12]. So $Z$ is non-singular if and only if $Z=Z_{\operatorname{pr}}$. Then Luna’s slice theorem implies that $Y\rightarrow Z$ is actually a $G$-fiber bundle with fiber $G/H$ (see [BKS25, Proposition 6.4]), where $(H)$ is the principal isotropy class.
@@
Let $Y$ be an irreducible non-singular affine $G$-variety with FIPG. Assume that the identity component $G^{0}$ is a torus. By [Sch95, Proposition 10.1], $Y$ is $k$-large if and only if $Y$ is $k$-principal. If $G$ acts on $Y$ freely outside a locus of codimension at least two, then $Y$ is $2$-large. Theorem 1.1 implies that $T^{*}Y/\!\!/\!\!/G$ is a symplectic variety with terminal singularities that is the affine closure of $T^{*}(Y/\!/G)_{\operatorname{pr}}$. One particular example is that when $G$ is a finite group that acts freely on $Y^{\diamond}\subset Y$ with $\operatorname{codim}Y\setminus Y^{\diamond}\geq 2$, then $T^{*}Y/G$ is the affine closure of $T^{*}(Y^{\diamond}/G)$, as previously remarked in [FL25a, Example 2.8].
@@ 4.2. Factoriality
For an irreducible non-singular affine $G$-variety $Y$, we remark that it follows from the standard properties of affine GIT that the natural projections $Y\rightarrow Z$ and $Y_{\operatorname{pr}}\rightarrow Z_{\operatorname{pr}}$ are both quotients by $G$ in the sense of [KKV89, 4.1].
@@
Since $Y$ is $2$-principal, it follows that $\operatorname{codim}(Y\setminus Y_{\operatorname{pr}})\geq 2$ for the $G$-stable open subset $Y_{\operatorname{pr}}$ of $Y$. In particular, one gets
@@
Moreover, since $Z\setminus Z_{\operatorname{pr}}$ has codimension at least two in $Z$, $Z_{\operatorname{pr}}\subset Z_{\operatorname{reg}}$ and $Z$ is normal, we have $\operatorname{Cl}(Z)\cong\operatorname{Pic}(Z_{\operatorname{pr}})$. By [KKV89, Lemma 2.2], there exists an exact sequence
@@
Then applying [KKV89, Proposition 4.2] to the quotient $Y_{\operatorname{pr}}\rightarrow Z_{\operatorname{pr}}$ shows that $\operatorname{Pic}(Z_{\operatorname{pr}})$ is a subgroup of $\operatorname{Pic}_{G}(Y_{\operatorname{pr}})$, so the result follows. ∎
@@
implies that $H^{1}_{\operatorname{alg}}(G,\mathcal{O}(Y)^{\times})$ is trivial (resp. finite) if the character group $\mathfrak{X}(G)$ of $G$ is trivial (resp. finite) and hence if $G_{\operatorname{ab}}$ is trivial (resp. finite). ∎
Proof of Theorem 1.4.
Recall that the singular locus of terminal singularities has codimension at least three and it follows from [Fu03, Corollary 1.3] that a locally $\mathbb{Q}$-factorial singular symplectic variety does not admit any symplectic resolution if its singular locus is not of pure codimension two.
Since $G$ is semisimple by assumption, it follows that $G$ is connected and $G_{\operatorname{ab}}$ is trivial, therefore $T^{*}Y/\!\!/\!\!/G$ is a $\mathbb{Q}$-factorial symplectic variety with terminal singularities by Theorems 1.1 and 1.3. In particular, $T^{*}Y/\!\!/\!\!/G$ admits a symplectic resolution if and only if $T^{*}Y/\!\!/\!\!/G$ itself is non-singular, and hence if and only if Y//GY/\!/G is non-singular by Theorem 1.1. ∎
@@
Let $Y$ be an irreducible non-singular affine $G$-variety that is $2$-large. Assume that $\mathcal{O}(Y)^{\times}=\mathbb{C}^{\times}$, $Y^{G}\not$$\emptyset$ and $\operatorname{Pic}(Y)=0$.
@@
    The divisor class group $\operatorname{Cl}(Z)$ is isomorphic to the kernel of $\mathfrak{X}(G)\rightarrow\mathfrak{X}(H)$, where $(H)$ is the principal isotropy class.
@@
Applying [KKV89, Corollary 5.3] to the quotient $Y\rightarrow Z$ yields $\operatorname{Pic}(Z)=0$. Let $p_{Z}\colon M\rightarrow Z$ be the natural projection and let $\mathbf{Z}$ be the zero section of $p_{Z}$ with the isomorphism $\iota=p_{Z}|_{\mathbf{Z}}\colon\mathbf{Z}\longrightarrow Z$. Consider the following commutative diagram:
@@
By Theorem 1.3 and Proposition 4.7, we only need to prove that if $M$ is $\mathbb{Q}$-factorial, then $G_{\operatorname{ab}}$ is finite. Assume to the contrary that $G_{\operatorname{ab}}$ is not finite. Then we can choose a surjective character $\chi\colon G\rightarrow\mathbb{C}^{\times}$. In particular, $\chi^{n}\not=1$ for all $n\geq 1$. Nevertheless, since the principal isotropy gruop $H$ is finite, there exists $m\in\mathbb{Z}_{>0}$ such that $\chi^{m}|_{H}=1$, which then implies that $\operatorname{Cl}(Z)$ is not torsion and hence $M$ is not $\mathbb{Q}$-factorial by Proposition 4.7, which is absurd. ∎
The following result generalizes [BS19, Corollary 1.3].
Corollary 4.9.
Under the notation and assumptions of Proposition 4.7, if $G_{\operatorname{ab}}$ is finite and $G$ acts non-trivially on $Y$, then $M$ does not admit a symplectic resolution.
Proof.
By Theorem 1.1 and Theorem 1.3, it remains to show that $Z$ is singular. Indeed, by [Sch95, Theorem 9.12], we have $Z_{\operatorname{reg}}=Z_{\operatorname{pr}}$. In particular, if $Z$ is non-singular, then $Z_{\operatorname{pr}}=Z$ and $Y_{\operatorname{pr}}=Y$, and it follows that $G$ is a finite group acting trivially on $Y$ as $Y^{G}\not$$\emptyset$, which is a contradiction. ∎
@@ 4.3. Examples
We gather some examples to illustrate the sharpness of Theorems 1.1 and 1.3.
@@
Let $V=(\mathfrak{sl}_{2})^{2}$, as a representation of $G=\operatorname{PGL}_{2}$, which is $1$-large, but not $2$-large. The Hamiltonian reduction $T^{*}V/\!\!/\!\!/G$ is identified with the locus of square zero matrices in $\mathfrak{sp}_{4}$, and in particular is a symplectic variety. Nevertheless, its singular locus is the codimension two locus consisting of rank one matrices in $\mathfrak{sp}_{4}$ and therefore it is not terminal, so the part (1.1.1) of Theorem 1.1 is optimal.
@@
It is easy to see that for any $k$, the Hamiltonian reduction $T^{*}V_{k}/\!\!/\!\!/G$ is isomorphic to the minimal nilpotent orbit closure $\overline{\mathbf{O}}_{{\rm min}}$ in $\mathfrak{sl}_{n}$. For $z\in V_{k}$, the $G$-orbit GzGz is closed in $V_{k}$ if and only if $z=0$ or $z\in V_{k}\setminus(F_{k}\cup F^{\prime}_{n-k})$, where $F_{k}=\{x_{1}=\cdots=x_{k}=0\}$ and $F^{\prime}_{n-k}=\{x_{k+1}=\cdots=x_{n}=0\}$. It follows that $(V_{k})_{\operatorname{pr}}=V_{k}\setminus(F_{k}\cup F^{\prime}_{n-k})$, so $V_{k}$ is 2-large if and only if $2\leq k\leq n-2$. In this case, Theorem 1.1 implies that $\overline{\mathbf{O}}_{{\rm min}}$ is the affine closure of $T^{*}((V_{k}/\!/G)_{\operatorname{pr}})$. On the other hand, $(V_{k}/\!/G)_{\operatorname{pr}}$ is just the punctured affine cone of the Segre variety $\mathbb{P}^{k-1}\times\mathbb{P}^{n-k-1}\subset\mathbb{P}^{k(n-k)-1}$. This shows that $\overline{\mathbf{O}}_{{\rm min}}$ is the affine closure of the cotangent bundle of the punctured affine cone of a Segre variety, which is a special case of [FL25a, Theorem 1.4] but with a new proof. For $k=0$, $V_{0}/\!/G$ is just one point. For $k=1$, $V_{1}$ is $1$-large with $V_{1}/\!/G\cong\mathbb{C}^{n-1}$ and $(V_{1}/\!/G)_{\operatorname{pr}}\cong\mathbb{C}^{n-1}\setminus\{0\}$. It follows that the affine closure of $T^{*}(V_{1}/\!/G)_{\operatorname{pr}}$ is $T^{*}\mathbb{C}^{n-1}$ which is different from $\overline{\mathbf{O}}_{{\rm min}}$. Hence, parts (1.1.2) and (1.1.3) of Theorem 1.1 are sharp.
@@
Following the notation in Example 4.11, we have $G_{\operatorname{ab}}=G=\mathbb{C}^{\times}$. Moreover, it is well-known that $\overline{\mathbf{O}}_{{\rm min}}$ is not $\mathbb{Q}$-factorial and admits a symplectic resolution by collapsing the zero section of $T^{*}\mathbb{P}^{n-1}$.
Now, let $n=2k$ and let $G^{\prime}\coloneqq\mathbb{C}^{\times}\rtimes\mathbb{Z}_{2}$, where $s\in\mathbb{Z}_{2}$ acts on $\mathbb{C}^{\times}$ by $s(\lambda)=\lambda^{-1}$. Then $G^{\prime},G^{\prime}]=\mathbb{C}^{\times}$ and $G_{\operatorname{ab}}\cong\mathbb{Z}_{2}$. Let $s$ act on $V_{k}=\mathbb{C}^{2k}$ as follows:
@@
The nilpotent closure $\overline{\textbf{O}}_{[2^{2},1^{2k-4}]}$ is next to the minimal orbit closure in the sense that it is contained in all other nilpotent orbit closures, and $\operatorname{Cl}(\overline{\textbf{O}}_{[2^{2},1^{2k-4}]})\cong\mathbb{Z}_{2}$ by [Fu03, Corollary 2.7]. These examples show that we cannot remove the triviality and finiteness conditions on $G_{\operatorname{ab}}$ in Theorem 1.3. In this sense, our Theorem 1.3 is also optimal.